# HA-VLN: A Benchmark for Human-Aware Navigation in Discrete–Continuous Environments with Dynamic Multi-Human Interactions, Real-World Validation, and an Open Leaderboard

### v1

@@ B.8 Human Activities Annotation Data Analysis
Human Distribution by Region. Figure A5(a) illustrates how 910 human figures are dispersed across 90 buildings, averaging about nine individuals per building. Even though each person moves independently, this distribution ensures robust and dynamic multi-human interactions, closely mirroring real-world scenarios.
Motion Frequency Analysis. Figures A5(b)–(c) compare the 15 most frequent motions before and after multi-human enrichments. While the total number of motions increases, we also embed additional movement patterns and group interactions into existing actions. For instance, “talking on the phone while pacing” may now involve extended pacing distances or layered scenarios like “a couple having a quiet conversation” or “friends sharing stories over dinner.”
Movement Distance Analysis. Figure A5(d) displays the distribution of trajectory lengths for actively moving humans. Specifically, 22.4% cover distances up to 0.5 m, 37.3% reach 0.5–1 m, 25.0% span 1–1.5 m, 11.6% extend 1.5–2 m, and the remaining 3.7% exceed 2 m. This wide range reflects the diverse indoor and outdoor behaviors encompassed in the dataset.
Human Impact Analysis. As shown in Figure A5(e), humans exert a notable influence on navigation paths: 35.5% of the 16,844 paths in HA-VLN physically intersect with human motion, while 49.7% of viewpoints are indirectly affected (i.e., humans are visible along the route). These statistics underline the importance of accounting for human presence and movement trajectories when designing real-world navigation agents.
@@ C.2 HA-R2R Instruction Generation
To create enriched instructions for HA-R2R, we use ChatGPT-4o and LLaMA-3-8B-Instruct to expand upon R2R-CE’s original textual data. Our strategy involves a carefully crafted few-shot prompt, combining a system prompt (Listing LABEL:list3:enrichment_prompt) and few-shot examples (Listing LABEL:list4:enrichment-few-shot-example).
Prompt Structure. The system prompt lays out guidelines for generating instructions that emphasize social context. It encourages mentioning human activities and interactions relevant to navigation paths. Few-shot examples then illustrate the desired format, including references to human behavior (e.g., “someone quietly making a phone call; keep your voice down as you proceed”), positional references, and object interactions.
@@ C.3 HA-R2R Data Analysis
Word Frequency Analysis. We conduct a word frequency study on HA-R2R to gauge its capacity for representing realistic, human-centric scenarios. Figures A6(a) and (b) illustrate frequently used nouns and verbs, respectively, confirming the dataset’s focus on both spatial navigation and social interactions.
@@
Human Impact Analysis. Figure A6(c) shows that most instructions contain 20–60% human-related content, reflecting the dataset’s emphasis on people in everyday scenes.
Comparisons of word clouds in Figures A6(d) and (e) confirm that while both human-aligned and non-human segments use common navigational verbs (walk, left, right), instructions involving humans introduce additional social context (couple, man, painting). This integration of interpersonal cues elevates HA-R2R beyond simple route directives, better mirroring real-world navigation challenges in human-filled environments.
@@ C.4 Visual and Depth Embeddings
Following VLN-CE [28], we employ parallel streams to process RGB and depth images. Each viewpoint produces a set of features from two specialized ResNet-50 models:
@@
  • Depth Features. Let $\{v^{d}_{1},v^{d}_{2},\dots,v^{d}_{k}\}$, where $v^{d}_{i}\in\mathbb{R}^{128}$, be outputs of another ResNet-50 pretrained on Gibson-4+ [53] and MP3D for point-goal navigation.
@@
$v_{i}=\bigl{[}\;v^{rgb}_{i}\,W_{\mathrm{rgb}};\;\,v^{d}_{i}\,W_{\mathrm{depth}};\;\,d_{i}\;\bigr{]}\,W_{\mathrm{merge}},$
@@ C.5 Text Embeddings
For the HA-VLN-VL agent, we utilize text embeddings from PREVALENT [18], which was pre-trained on 6.58M image–text–action triplets, thereby capturing broad contextual cues for navigation. Conversely, the HA-VLN-CMA agent adopts embeddings from BERT [10], also widely used for its strong language representations.
Formally, let =$\ell=\{w_{1},\dots,w_{n}\}$ be a sequence of tokens representing the instruction. Each token $w_{i}$ is mapped to a one-hot vector $e_{i}\in\mathbb{R}^{V}$, where $V$ is the vocabulary size. An embedding matrix $E\in\mathbb{R}^{V\times d}$ then projects $e_{i}$ into a continuous $d$-dimensional space:
$x_{i}=E^{\top}\,e_{i},\quad x_{i}\in\mathbb{R}^{d}.$
@@
Figure A8: Trajectory Comparison Under Human vs. No-Human Conditions. We illustrate the same episode’s trajectories predicted by BEVBert [1], trained on VLN-CE, in scenarios with (bottom) and without (top) human presence. In the top row, no bystanders are present, and the agent follows its instructions with minimal collision risk. In the bottom row, bystanders and human-aligned cues lead to altered motion decisions, sometimes creating additional collision challenges or deviations.
Figure A9: Agent Trajectory Examples (HA-VLN-CMA). The top row depicts a failed navigation case in which the agent fails to avoid an oncoming human, resulting in a collision. By contrast, the bottom row illustrates a successful run: the agent proactively adjusts its path to the right, thereby avoiding interference and completing the task collision-free.
@@ C.6 HA-VLN-VL Structure
Model Overview. HA-VLN-VL adopts a BERT-like architecture inspired by Recurrent VLN-BERT [20], extending it to handle human-aware navigation. At each timestep $t$, the model receives the previous state $s_{t-1}$, language tokens $X$, and fused RGB–depth visual features $V_{t}$ (Sec. C.4). It outputs an updated state $s_{t}$ and an action distribution $p_{t}^{a}$:
$s_{t},\;p_{t}^{a}=\mathrm{HA\text{-}VLN\text{-}VL}\bigl{(}s_{t-1},\,X,\,V_{t}\bigr{)}.$
State Token. In line with BERT conventions, the model maintains a state token $s_{t}$ that encapsulates the agent’s internal context. Initially, $s_{0}$ is set to the embedding of the [CLS] token. At each step, the state token is updated by appending the agent’s previously executed action $a_{t}$ and projecting the resulting vector:
$s_{t}=[\,s^{\prime}_{t};\,a_{t}\,]\;W_{s},$
@@
$A_{s,v}^{t}=\mathrm{Softmax}\!\Bigl{(}\frac{Q_{s}\,K_{v}^{\top}}{\sqrt{d_{h}}}\Bigr{)},$
@@
$p_{t}^{a}=\mathrm{\overline{AveragePool}}\bigl{(}A_{s,v^{\prime}}^{t}\bigr{)},$
@@
$a_{t}=\arg\max\bigl{(}p_{t}^{a}\bigr{)}.$
Training Objective. HA-VLN-VL is optimized through a combination of supervised imitation learning—to mimic ground-truth trajectories—and optional reinforcement learning, which rewards safe and efficient paths. As depicted in Figure A7(a), the model continuously refines its understanding of language instructions and visual cues, offering robust and socially aware navigation.
@@ C.7 HA-VLN-CMA Structure
Architecture Overview. HA-VLN-CMA is a dual-stream visual-language agent featuring Cross-Modal Attention (CMA) and a recurrent decoder for navigation in human-populated scenarios (see Figure A7(b)). It processes two visual channels—RGB and Depth—alongside language instructions, then outputs an action at each time step.
@@
$v^{\mathrm{rgb}}_{t}=\mathrm{ResNet}^{\mathrm{rgb}}(o_{t}),\quad v^{\mathrm{d}}_{t}=\mathrm{ResNet}^{\mathrm{depth}}(o_{t}),$
@@
$v_{i}=\bigl{[}v^{\mathrm{rgb}}_{i}W_{\mathrm{rgb}};\;v^{d}_{i}W_{\mathrm{depth}};\;d_{i}\bigr{]}\,W_{\mathrm{merge}},$
@@
$l=\mathrm{BERT}(w_{1},\dots,w_{T}).$
@@
$f_{t}=\mathrm{GRU}\bigl{(}\,[\,(\,v_{t},l),\,a_{t-1}\,],\,f_{t-1}\bigr{)},$
@@
$a_{t}=\mathrm{softmax}(\mathrm{MLP}(f_{t})),$
@@ D.1 Evaluation Metrics
Navigation Error (NE). NE is the mean distance between the agent’s final position and the intended target:
$\mathrm{NE}=\frac{\sum_{i=1}^{L}d_{i}}{L},$
@@
Success Rate (SR). SR measures the ratio of episodes completed with zero collisions:
$\mathrm{SR}=\frac{\sum_{i=1}^{L}\mathbb{I}\bigl{(}c_{i}-|A^{c}_{i}|=0\bigr{)}}{L},$
@@ D.2 Ground Truth Path Annotation
D.3 Visualization of Navigation
Figures A8A9 illustrate trajectories predicted by BEVBert [1] (trained on VLN-CE) and HA-VLN-CMA, which showcases success and failure in human-filled or empty environments.
Failures with Human Crossing. In Figure A8, the agent performs well when no bystanders are present. Yet in a human-populated setting, it fails to adjust at step 37 when a volunteer crosses its path, leading to collision.
Collision vs. Avoidance. Figure A9 similarly shows two scenarios. At step 39 in the top pane, a direct approach by a bystander overwhelms the agent, causing a collision. In the bottom pane at step 22, the agent successfully deviates upon sensing a person nearby, avoiding any collision altogether. These visualizations confirm that dynamic human presence greatly complicates navigation, highlighting the need for robust social-aware models.
Figure A10: Unitree GO2-EDU quadruped robot used in real-world experiments. The platform integrates an NVIDIA Jetson Orin NX for onboard computation, an IMU for pose estimation, an Intel RealSense D435i camera for visual perception, and a 3D LiDAR for obstacle/human detection.
Figure A11: Navigation success in an office setting (left: no humans, right: with humans). Top: The given instruction for the robot. Middle: A third-person view of the robot’s path. Bottom: The robot’s egocentric camera feed.
Figure A12: Navigation failure in an office setting. A volunteer abruptly changes position, causing the robot to collide mid-path. This scenario highlights the difficulty of adapting to sudden human movement in confined workspaces.
@@ D.3 Visualization of Navigation
Figure A14: Navigation failure in a living room with multiple bystanders. Attempting to move beyond the sofa toward a dining area, the robot collides with a volunteer who abruptly stands and shifts position. This underscores how unpredictable human motion can disrupt the agent’s intended path, requiring rapid re-planning.
@@
D.4 Validation on Real-World Robots
@@
Robot Setup. Our evaluations use a Unitree GO2-EDU quadruped (Figure A10), featuring an Intel Realsense D435i camera providing RGB imagery at a $90^{\circ}$ field of view and a 3D LiDAR below the camera for obstacle detection. The IMU refines positional and orientational control, enabling consistent motion execution.
Visual Demonstrations. Figures A11, A13, and A15 show the robot traversing distinct indoor environments—offices, living rooms, and hallways—guided by natural-language instructions. In Figure 5, the robot navigates around multiple people, leveraging camera inputs to avoid collisions through minor path adjustments. Although the agent typically succeeds in reaching its destination, collisions remain possible when bystanders change their trajectories unexpectedly. Figures A12, A14, and A16 illustrate such scenarios, highlighting real-time challenges in unpredictable, human-inhabited spaces. Additional demonstrations, including a compilation video on our project webpage, further illustrate the robot’s performance and underscore how human-aware training significantly aids sim-to-real transfer in dynamic indoor environments.

### v3

@@ B.8 Human Activities Annotation Data Analysis
Human Distribution by Region. Fig. A4(a) illustrates the distribution of 910 humans across 26 region types in 90 buildings, averaging about nine individuals per building. Even though each person moves independently, this distribution ensures robust and dynamic multi-human interactions, closely mirroring real-world scenarios.
Motion Frequency Analysis. Figs. A4(b)–(c) compare the 15 most frequent motions before and after multi-human enrichments. While the total number of motions increases, we also embed additional movement patterns and group interactions into existing actions. For instance, “talking on the phone while pacing” may now involve extended pacing distances or layered scenarios like “a couple having a quiet conversation” or “friends sharing stories over dinner.”
Movement Distance Analysis. Fig. A4(d) displays the distribution of trajectory lengths for actively moving humans. Specifically, 22.4% cover distances up to 0.5 m, 37.3% reach 0.5–1 m, 25.0% span 1–1.5 m, 11.6% extend 1.5–2 m, and the remaining 3.7% exceed 2 m. This wide range reflects the diverse indoor and outdoor behaviors encompassed in the dataset.
Human Impact Analysis. As shown in Fig. A4(e), humans exert a notable influence on navigation paths: 35.5% of the 16,844 paths in HA-VLN physically intersect with human motion, while 49.7% of viewpoints are indirectly affected (i.e., humans are visible along the route). These statistics underline the importance of accounting for human presence and movement trajectories when designing real-world navigation agents.
@@ C.2 HA-R2R Instruction Generation
To create enriched instructions for HA-R2R, we use ChatGPT-4o and LLaMA-3-8B-Instruct to expand upon R2R-CE’s original textual data. Our strategy involves a carefully crafted few-shot prompt, combining a system prompt (Listing 3) and few-shot examples (Listing 4).
Prompt Structure. The system prompt lays out guidelines for generating instructions that emphasize social context. It encourages mentioning human activities and interactions relevant to navigation paths Wu et al. (2025). Few-shot examples then illustrate the desired format, including references to human behavior (e.g., “someone quietly making a phone call; keep your voice down as you proceed”), positional references, and object interactions.
@@ C.3 HA-R2R Data Analysis
Word Frequency Analysis. We conduct a word frequency study on HA-R2R to gauge its capacity for representing realistic, human-centric scenarios. Figs. A6(a) and (b) illustrate frequently used nouns and verbs, confirming the dataset’s focus on both spatial navigation and social interactions.
@@
Human Impact Analysis. Fig. A6(c) shows that most instructions contain 20–60% human-related content, reflecting the dataset’s emphasis on people in everyday scenes.
Comparisons of word clouds in Figs. A6(d) and (e) confirm that while both human-aligned and non-human segments use common navigational verbs (walk, left, right), instructions involving humans introduce additional social context (couple, man, painting). This integration of interpersonal cues elevates HA-R2R beyond simple route directives, better mirroring real-world navigation challenges in human-filled environments.
@@ C.4 Visual and Depth Embeddings
Following VLN-CE Krantz et al. (2020), we employ parallel streams to process RGB and depth images. Each viewpoint produces a set of features from two specialized ResNet-50 models:
  1. 1.
@@
  2. 2.
    Depth Features. Let $\{v^{d}_{1},v^{d}_{2},\dots,v^{d}_{k}\}$, where $v^{d}_{i}\in\mathbb{R}^{128}$, be outputs of another ResNet-50 pretrained on Gibson-4+ Xia et al. (2018) and MP3D for point-goal navigation.
@@
$v_{i}=\bigl[\;v^{rgb}_{i}\,W_{\mathrm{rgb}};\;\,v^{d}_{i}\,W_{\mathrm{depth}};\;\,d_{i}\;\bigr]\,W_{\mathrm{merge}},$ (A2)
@@ C.5 Text Embeddings
For the HA-VLN-VL agent, we utilize text embeddings from PREVALENT Hao et al. (2020), which was pre-trained on 6.58M image–text–action triplets, thereby capturing broad contextual cues for navigation. Conversely, the HA-VLN-CMA agent adopts embeddings from BERT Devlin (2018), also widely used for its strong language representations.
Formally, let $\ell=\{w_{1},\dots,w_{n}\}$ be a sequence of tokens representing the instruction. Each token $w_{i}$ is mapped to a one-hot vector $e_{i}\in\mathbb{R}^{V}$, where $V$ is the vocabulary size. An embedding matrix $E\in\mathbb{R}^{V\times d}$ then projects $e_{i}$ into a continuous $d$-dimensional space:
$x_{i}=E^{\top}\,e_{i},\quad x_{i}\in\mathbb{R}^{d}.$ (A3)
@@
Figure A8: Trajectory Comparison Under Human vs. No-Human Conditions. We illustrate the same episode’s trajectories predicted by BEVBert An et al. (2023), trained on VLN-CE, in scenarios with (bottom) and without (top) human presence. In the top row, no bystanders are present, and the agent follows its instructions with minimal collision risk. In the bottom row, bystanders and human-aligned cues lead to altered motion decisions, sometimes creating additional collision challenges or deviations.
@@ C.6 HA-VLN-VL Structure
Model Overview. HA-VLN-VL adopts a BERT-like architecture inspired by Recurrent VLN-BERT Hong et al. (2021), extending it to handle human-aware navigation. At each timestep $t$, the model receives the previous state $s_{t-1}$, language tokens $X$, and fused RGB–depth visual features $V_{t}$ (Sec. C.4). It outputs an updated state $s_{t}$ and an action distribution $p_{t}^{a}$:
$s_{t},\;p_{t}^{a}=\mathrm{HA\text{-}VLN\text{-}VL}\bigl(s_{t-1},\,X,\,V_{t}\bigr).$ (A4)
State Token. In line with BERT conventions, the model maintains a state token $s_{t}$ that encapsulates the agent’s internal context. Initially, $s_{0}$ is set to embedding of [CLS] token. At each step, the state token is updated by appending agent’s previously executed action $a_{t}$ and projecting resulting vector:
$s_{t}=[\,s^{\prime}_{t};\,a_{t}\,]\;W_{s},$ (A5)
@@
$A_{s,v}^{t}=\mathrm{Softmax}\!\Bigl(\frac{Q_{s}\,K_{v}^{\top}}{\sqrt{d_{h}}}\Bigr),$ (A6)
@@
$p_{t}^{a}=\mathrm{\overline{AveragePool}}\bigl(A_{s,v^{\prime}}^{t}\bigr),$ (A7)
@@
$a_{t}=\arg\max\bigl(p_{t}^{a}\bigr).$ (A8)
Training Objective. HA-VLN-VL is optimized through a combination of supervised imitation learning—to mimic ground-truth trajectories—and optional reinforcement learning, which rewards safe and efficient paths. As depicted in Fig. A7(a), the model continuously refines its understanding of language instructions and visual cues, offering robust and socially aware navigation.
@@ C.7 HA-VLN-CMA Structure
Architecture Overview. HA-VLN-CMA is a dual-stream visual-language agent featuring Cross-Modal Attention (CMA) and a recurrent decoder for navigation in human-populated scenarios (see Fig. A7(b)). It processes two visual channels—RGB and Depth—alongside language instructions, then outputs an action at each time step.
@@
$v^{\mathrm{rgb}}_{t}=\mathrm{ResNet}^{\mathrm{rgb}}(o_{t}),\quad v^{\mathrm{d}}_{t}=\mathrm{ResNet}^{\mathrm{depth}}(o_{t}),$ (A9)
@@
$v_{i}=\bigl[v^{\mathrm{rgb}}_{i}W_{\mathrm{rgb}};\;v^{d}_{i}W_{\mathrm{depth}};\;d_{i}\bigr]\,W_{\mathrm{merge}},$ (A10)
@@
$l=\mathrm{BERT}(w_{1},\dots,w_{T}).$ (A11)
@@
$f_{t}=\mathrm{GRU}\bigl(\,[\,(\,v_{t},l),\,a_{t-1}\,],\,f_{t-1}\bigr),$ (A13)
@@
$a_{t}=\mathrm{softmax}(\mathrm{MLP}(f_{t})),$ (A14)
@@ D.1 Evaluation Metrics
Navigation Error (NE). NE is the mean distance between agent’s final position and intended target:
$\mathrm{NE}=\frac{\sum_{i=1}^{L}d_{i}}{L},$ (A17)
@@
Success Rate (SR). SR measures the ratio of episodes completed with zero collisions, and checks if the agent stops sufficiently close to the goal Anderson et al. (2018), we provide the equation for the collision check part here:
$\mathrm{SR}=\frac{\sum_{i=1}^{L}\mathbb{I}\bigl(c_{i}-|A^{c}_{i}|=0\bigr)}{L},$ (A18)
@@ D.2 Ground Truth Path Annotation
D.3 Further Discussion on Step Size
In Table A4, a 1.0 m step was treated as four 0.25 m sub-steps, and a 2.25 m step as nine 0.25 m sub-steps, with collisions checked after each sub-step. When evaluated on the val_unseen split, BEVBert agent fails to navigate effectively with both 1.0 m and 2.25 m step sizes (SR drops to zero).
Table A4: Impact of Step Size Combination on Navigation. In this experiment, we treat 1m step as four 0.25m steps, and 2.25m step as nine 0.25m steps. In this case, collisions are detected every 0.25m. We show results for BEVBert An et al. (2023) on unseen validation.
[…]
1.00 6.85 26.97 0.94 0.004
2.25 8.79 112.78 0.97 0.000
D.4 Visualization of Navigation
Figs. A85 illustrate trajectories predicted by BEVBert An et al. (2023) (trained on VLN-CE) and HA-VLN-CMA, which showcases success and failure in human-filled or empty environments.
Failures with Human Crossing. In Fig. A8, the agent performs well when no bystanders are present. Yet in a human-populated setting, it fails to adjust at step 37 when a volunteer crosses its path, leading to collision.
Collision vs. Avoidance. Fig. 5 similarly shows two scenarios. At step 39 in the top pane, a direct approach by a bystander overwhelms the agent, causing a collision. In the bottom pane at step 22, the agent successfully deviates upon sensing a person nearby, avoiding any collision altogether. These visualizations confirm that dynamic human presence greatly complicates navigation, highlighting the need for robust social-aware models.
Figure A9: Navigation success in an office (left: no humans, right: with humans). Top: The given instruction for the robot. Middle: A third-person view of the robot’s path. Bottom: The robot’s selected view.
Figure A10: Navigation failure in an office setting. A volunteer abruptly changes position, causing robot to collide mid-path. This highlights the difficulty of adapting to sudden human movement in confined workspaces.
@@ D.4 Visualization of Navigation
Figure A12: Navigation failure in a living room with multiple bystanders. Attempting to move beyond sofa toward a dining area, the robot collides with a volunteer who abruptly stands and shifts position. This underscores how unpredictable human motion can disrupt agent’s intended path, requiring rapid re-planning.
@@
D.5 Validation on Real-World Robots
@@
Table A5: Navigation success rate across different region layouts with (w/) and without (w/o) human presence. Each result is averaged over 30 episodes across 3 instances of each region type.
Living Room Office Hallway Lobby ALL
Methods w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
HA-VLN-CMA-Base (trained on VLN-CE) 0.23 0.08 0.26 0.08 0.30 0.13 0.24 0.07 0.26 0.09
HA-VLN-VL (trained on VLN-CE) 0.38 0.11 0.38 0.10 0.47 0.17 0.38 0.10 0.40 0.12
HA-VLN-CMA-Base (trained on HA-VLN) 0.24 0.13 0.24 0.13 0.29 0.20 0.23 0.13 0.25 0.15
HA-VLN-VL (trained on HA-VLN) 0.42 0.17 0.43 0.17 0.49 0.20 0.43 0.17 0.44 0.18
Setup. Our evaluations use a Unitree GO2-EDU quadruped, featuring the Intel Realsense D435i camera providing RGB imagery and a 3D LiDAR below camera for detection. IMU refines positional and orientational control, enabling consistent motions. The quadruped rotates to get the panoramic view at each step. We evaluate our agents in four types of everyday indoor environments (each with three instances)—office, living room, hallway, and lobby—under two conditions: (i) w/o human presence (no bystanders) and (ii) w/ human presence (2-4 free-moving volunteers). This setup simulates realistic indoor traffic patterns and partial observability.
Observations. As illustrated in Fig. 6 (b), the robot frequently pauses or yields to avoid oncoming pedestrians. In the absence of bystanders, it navigates smoothly (Fig. A9), but collisions arise in cramped corridors or when crowds converge suddenly (Fig. A10). We observe similar patterns in living-room environments (Figs. A11A12) and hallways (Fig. A13).
Table A5 shows the average NSR (Navigation Success Rate) across 30 trials in each instance. While human presence invariably lowers NSR, HA-VLN-VL consistently outperforms HA-VLN-CMA-Base, demonstrating stronger adaptability to dynamic motion. Also, Table A5 shows agents trained on HA-VLN achieve higher NSR (0.18 vs. 0.12) than VLN-CE, demonstrating HA-R2R’s sim-to-real gain under realistic conditions. Still, partial observability and abrupt group formations remain challenging, especially in narrow passages or at congested junctions. Appendix D.5 further details performance under varying crowd densities.
Visual Demonstrations. Figs. A9, A11, and A13 show the robot traversing distinct indoor environments—offices, living rooms, and hallways—guided by natural-language instructions. In Fig. 6 (b), the robot navigates around multiple people, leveraging camera inputs to avoid collisions through minor path adjustments. Although the agent typically succeeds in reaching its destination, collisions remain possible when bystanders change their trajectories unexpectedly. Figs. A10, A12, and A14 illustrate such scenarios, highlighting real-time challenges in unpredictable, human-inhabited spaces. More demos on our project webpage, further illustrate robot’s performance and underscore how human-aware training aids sim-to-real transfer in dynamic indoor environments.
Insights. These experiments confirm that simulation-trained, multi-human navigation policies can indeed transfer to physical robots. However, further refinement in collision forecasting and reactive control is needed to handle unpredictable human behavior in tight indoor settings.
Appendix E Use of LLMs
Large Language Models (LLMs) were used to aid in the writing and polishing of the manuscript. Specifically, we used an LLM to assist in refining the language, improving readability, and ensuring clarity in various sections of the paper. The model helped with tasks such as sentence rephrasing, grammar checking, and enhancing the overall flow of the text.
It is important to note that the LLM was not involved in the ideation, research methodology, or experimental design. All research concepts, ideas, and analyses were developed and conducted by the authors. The contributions of the LLM were solely focused on improving the linguistic quality of the paper, with no involvement in the scientific content or data analysis.
The authors take full responsibility for the content of the manuscript, including any text generated or polished by the LLM. We have ensured that the LLM-generated text adheres to ethical guidelines and does not contribute to plagiarism or scientific misconduct.